A value commonly taken for the perception reaction time of motorists is delta t equals 1.5 seconds. Using this value determine the total stopping distance for a car traveling at 30 m/s. In calculating the stopping distance, lets assume that the vehicle continues at its initial speed for delta t, then decelerates to rest. A reasonable value for the deceleration in braking is 8.0 m/s^2
Answer #1

Okay, total distance comprises of two parts

1) reaction distance

2) Braking distance

reaction distance = speed * time

reaction distance = 30 * 1.5 = 45 m

Now, braking distance = v2 / 2a

braking distance = 900/ 2 * 8

braking distance = 56.25 m

so,

Total distance = 101.25 m